### v1

# envname-P envname#1

Abstract
Quantization-aware training (QAT) is a leading technique for improving the accuracy of quantized neural networks. Previous work has shown that decomposing training into a full-precision (FP) phase followed by a QAT phase yields superior accuracy compared to QAT alone. However, the optimal allocation of compute between the FP and QAT phases remains unclear. We conduct extensive experiments with various compute budgets, QAT bit widths, and model sizes from 86.0M to 2.2B to investigate how different QAT durations impact final performance. We demonstrate that, contrary to previous findings, the loss-optimal ratio of QAT to FP training increases with the total amount of compute. Moreover, the optimal fraction can be accurately predicted for a wide range of model sizes and quantization widths using the tokens-per-parameter-byte statistic. From experimental data, we derive a loss scaling law that predicts both optimal QAT ratios and final model performance across different QAT/FP compute allocation strategies and QAT bit widths. We use the scaling law to make further predictions, which we verify experimentally, including which QAT bit width is optimal under a given memory constraint and how QAT accuracy with different bit widths compares to full-precision model accuracy. Additionally, we propose a novel cooldown and QAT fusion approach that performs learning rate decay jointly with quantization-aware training, eliminating redundant full-precision model updates and achieving significant compute savings. These findings provide practical insights into efficient QAT planning and enable the training of higher-quality quantized models with the same compute budget.
[…]
[Correspondence] {alexdremov, grangier, a_katharopoulos, awni}@apple.com * Work done at Apple during internship while studying at École Polytechnique Fédérale de Lausanne (EPFL), aleksandr.dremov@epfl.ch.
1 Introduction
As Large Language Models (LLMs) grow in size and on-device applications gain traction (Wahab & Adda, 2025), significant attention has been devoted to reducing inference costs via model compression (Frantar et al., 2022; Lin et al., 2024; Ma et al., 2023). One state-of-the-art method is quantization-aware training (QAT) (Chen et al., 2025a; Lin et al., 2024; Liu et al., 2025; Jacob et al., 2018). To adapt the model to the loss of numerical precision, QAT incorporates quantization directly into the model training process. It has been shown that QAT outperforms post-training quantization (PTQ) (Xiao et al., 2023; Banner et al., 2019), where quantization is applied after training is completed. Moreover, Liu et al. (2025) demonstrated that the best accuracy is achieved when a QAT phase follows a full-precision (FP) training phase.
For models designed for on-device use, the QAT stage is an important part of the training process and is usually planned in advance. As model sizes grow and deployment constraints tighten, practitioners face a critical resource allocation problem: given a fixed compute budget, how should training time be divided between full-precision pretraining and quantization-aware training? This decision directly impacts both model quality and deployment efficiency, yet existing guidelines assume fixed allocation ratios regardless of scale. As motivation, we note that Kumar et al. (2025) demonstrated that the error introduced by post-training quantization grows with the size of the pretraining data, which can actually make additional pretraining harmful. Intuitively, analogous to PTQ, having a longer full-precision stage should make subsequent QAT more difficult. While Liu et al. (2025) showed that spending 10% of the training steps on QAT is optimal for some setups, the authors did not explore how this proportion varies across different training lengths and model sizes.
In this work, we show that previous conclusions about optimal QAT length do not hold with an increased compute budget. Through a series of experiments with different model sizes and token counts, we demonstrate that the optimal fraction of QAT compared to the total training length increases with the total compute budget. This optimum can be accurately predicted for a wide range of setups using the tokens-per-parameter-byte statistic. Additionally, we propose a loss scaling law as a function of model parameter count ($N$), token count spent on full-precision training ($D_{\text{fp}}$), token count spent on QAT ($D_{\text{qat}}$), and QAT bit-width ($B$). The fitted law accurately captures the growth of the optimal QAT fraction with compute scale. The key contributions of this study are:
  • Unlike previously assumed, we find that the optimal fraction allocated for QAT increases with the growth of the tokens-per-parameter-byte statistic. This finding allows higher-quality quantized models to be achieved with the same initial compute budget (figure 1 (Left)).
  • We propose a loss scaling law that captures the optimal QAT fraction phenomenon and models the final expected loss of the FP and QAT pipeline (figure 1 (Right)). We use the scaling law to make further predictions, including which QAT bit-width is optimal under a given memory constraint and how QAT accuracy with different bit-widths compares to full-precision model accuracy.
  • We propose a novel approach: cooldown & QAT fusion—a scheme where learning rate decay is performed jointly with quantization-aware training, eliminating redundant full-precision updates and achieving better accuracy for the same token count.
[…]
[width=]svg-inkscape/figure1_svg-tex.pdf_tex
Figure 1: On the left, experimental and predicted optimal QAT fractions as a function of tokens-per-parameter-byte are shown. Different colors represent models of varying sizes, while point sizes indicate final perplexity normalized across experiments with identical total token counts for each model size. Results span multiple QAT bit-widths, and optimal QAT fraction values for endpoints are displayed. The plot demonstrates that the optimal QAT fraction increases with the full training tokens-per-parameter-byte statistic. On the right, loss scaling law predictions for a 4-bit QAT 396M parameter model across varying QAT and FP training lengths. Both experimental and theoretical optima are shown. The optimal QAT fraction predicted by the loss scaling law for each total token count closely matches the experimentally observed fraction.
2 Related Work
Quantization of LLMs.
Quantization is a method for reducing both the memory footprint and the computational requirements of neural networks by lowering the precision used in the network. By reducing the bit-width of the weights, activations, or both, quantization enables models to run faster, consume less power, and use less memory, which is particularly beneficial for deployment on resource-constrained devices. There are different quantization techniques, including post-training quantization (PTQ) (Xiao et al., 2023; Banner et al., 2019) — methods that transform a model after training has been completed and usually require minimal additional computational usage. Another group of methods is quantization-aware training (QAT) (Chen et al., 2025a; Lin et al., 2024; Liu et al., 2025; Jacob et al., 2018), which quantizes a model during training, allowing the model to better adapt to precision loss. As quantization operations are non-differentiable, training relies on gradient approximations such as the straight-through estimator (Bengio et al., 2013). In contrast to PTQ, QAT requires more computation, as effectively the full model is trained with added quantization-related operations. In this work, we focus on QAT, as it is the method most commonly used in practice to obtain high-quality quantized models.
Loss Scaling Laws.
Multiple works have previously addressed the problem of predicting final model loss ($L$) as a function of parameter count ($N$) and consumed tokens ($D$) (Bi et al., 2024; Hoffmann et al., 2022b; Kaplan et al., 2020). The Chinchilla (Hoffmann et al., 2022b) loss model is one of the most commonly used: $L(N,D)=E+AN^{-\alpha}+CD^{-\beta},$ where $A$, $C$, $\alpha$, $\beta$, and $E$ are fitted parameters. Bi et al. (2024) expand on this idea, fitting accuracy as a function of used non-embedding FLOPs (FLOP estimation of model inference without embedding layer calculations), showing that such an approach works better across different model sizes. Additionally, they show that scaling laws are greatly influenced by data quality.
QAT Loss Scaling Laws.
Chen et al. (2025b) proposed scaling law modeling specifically for QAT loss, adding a QAT-related penalty to the Chinchilla loss model:
$L(N,D,G)=\underbrace{E+\frac{A}{N^{\alpha}}+\frac{C}{D^{\beta}}}_{\text{Chinchilla loss}}+\underbrace{\frac{k\cdot D^{\gamma_{D}}\cdot(\log_{2}{G})^{\gamma_{G}}}{N^{\gamma_{N}}}}_{\text{QAT error}},$ (2.1)
where $G$ is the quantization granularity (number of elements in each quantization group), $k$, $\gamma_{D}$, $\gamma_{G}$, and $\gamma_{N}$ are fitted parameters, and the Chinchilla loss parameters are fixed from the non-quantized model fit. Therefore, this approach effectively models QAT error relative to the full-precision model for the same token and parameter count. However, formulas are fitted exclusively for each quantization bit width, which complicates analysis of the relationship between different bit widths. Kumar et al. (2025) propose precision-aware scaling laws for training and inference, predicting loss from low-precision training and PTQ or QAT.
While the Kumar et al. (2025); Chen et al. (2025b) laws are useful for understanding the final accuracy of a model trained with QAT from scratch, they overlook the fact that QAT is typically resumed from full-precision training to achieve better accuracy (Liu et al., 2025; Zhou et al., 2025a). Our work addresses this issue and presents a novel loss scaling law that explicitly handles the case when QAT is started from a full-precision checkpoint and works across different bit widths.
3 Experimental Setup
We use a decoder-only transformer (Zhang et al., 2021) identical to Llama 2 (Touvron et al., 2023). The architecture incorporates SwiGLU activations (Shazeer, 2020), RoPE (Su et al., 2024), RMSNorm (Zhang & Sennrich, 2019), alternating attention and feed-forward layers, and tied embedding and language-modeling head weights. We use the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99,\varepsilon=10^{-8}$) with decoupled weight decay of $0.01$ (Loshchilov & Hutter, 2017; Dereich et al., 2024) for all parameters outside the embedding and normalization layers. All experiments are trained with bfloat16 automatic mixed precision (Liu et al., 2021). Training is conducted on the DCLM dataset (Li et al., 2024), tokenized with the Llama 2 tokenizer with a 32,000-token vocabulary. We merge all tokenized documents into a single sequence with appropriate delimiting tokens and take chunks of 1024 tokens (used sequence length) for the batch—an approach also known as “concat-and-chunk” (Pouransari et al., 2024). The dataset is split into training and validation sets, and validation perplexity is used for evaluation. For QAT algorithms, we rely on ParetoQ (Liu et al., 2025) for our setups, as this method achieves state-of-the-art accuracy across different bit widths by combining different approaches. Additional information on the training procedure is provided in appendix 9.
4 Optimal QAT Compute
To study how loss changes for different combinations of QAT/FP training length, we train models of different sizes, different FP stage token counts ($D_{\text{fp}}$), QAT token counts ($D_{\text{qat}}$), and different QAT bit widths $B$. For the smallest model (86.0M parameters), we conduct experiments from 2.3B to 1.4T total tokens, while for the largest model (759.0M parameters), we conduct experiments from 8.5B to 669.2B total tokens. We focus primarily on 1-, 2-, 4-, and 6-bit quantization widths. Token counts and QAT fractions for our experiments are reported in appendix 21. We also verify that the obtained post-QAT models maintain reasonable accuracy and present a comparison to full-precision models in appendix 22. Specifically, our 4- and 6-bit setups achieve quality close to that of the full-precision model for the same total token count, and the drop in quality for 1- and 2-bit is reasonable.
The main objective of this study is to determine the optimal QAT fraction $f^{*}$—the fraction of the token count that should be dedicated to QAT for a given total token count. This can be formalized as the following minimization problem:
$D_{\text{qat}}^{*}(N,D_{\text{total}},B)=\operatorname*{arg\,min}_{\begin{subarray}{c}D_{\text{qat}}\in{\mathbb{N}},\\ D_{\text{qat}}+D_{\text{fp}}=D_{\text{total}}\end{subarray}}L(N,D_{\text{fp}},D_{\text{qat}},B),\;\;\;f^{*}(N,D_{\text{total}},B)=\frac{D_{\text{qat}}^{*}(N,D_{\text{total}},B)}{D_{\text{total}}},$
[…]
[width=]svg-inkscape/tok_v_tpb_svg-tex.pdf_tex
Figure 2: On the top, QAT optima for 396M model plotted in token coordinates. Different optima for the same total token count and different QAT bit widths can be observed. On the bottom, QAT optima for 396M model plotted in tokens-per-parameter-byte coordinates. With byte adjustment, different bit widths lie on the proposed fit line better.
where $L(N,D_{\text{fp}},D_{\text{qat}},B)$ is the final loss of the setup with $D_{\text{fp}}$ tokens dedicated to FP training and $D_{\text{qat}}$ tokens dedicated to $B$-bit QAT. Intuitively, $f^{*}$ expresses a trade-off. On one hand, too few QAT steps do not allow the quantized model to adapt to reduced precision. On the other hand, too many QAT steps (at the expense of fewer FP steps) should also lead to worse loss since QAT is trained with gradient approximations for quantization operators, which introduces biased and noisier gradients. Naturally, a trade-off emerges, suggesting that such $f^{*}$ should be well-defined.
4.1 Predicting the Optimal QAT Fraction
In this section, we focus on fitting the optimal QAT fractions directly. To account for different QAT bit widths used, we introduce the tokens-per-parameter-byte statistic. This choice was made based on several observations: larger models are generally easier to quantize, models trained for longer are harder to quantize, and smaller QAT bit widths are harder to quantize as well. While being intuitive from a QAT accuracy perspective, it can also predict the optimal fraction with high precision. Figure 2 provides a comparison between the two approaches. It is clearly seen that using tokens-per-parameter-byte provides an interpretable adjustment, facilitating a better fit. For optimal QAT fraction prediction, we fit a function of the form
$\widehat{f}(D_{\text{total}},N,B)$ $=\frac{\exp\left(\log{S_{\text{total}}}-\frac{a}{\log{S_{\text{total}}}}\right)}{S_{\text{total}}},$
$S_{\text{total}}$ $=\frac{D_{\text{total}}}{N\cdot\frac{B}{8}},$
where $\widehat{f}$ is the predicted optimal QAT fraction for the total tokens-per-parameter-byte count $S_{\text{total}}$, and $a$ is a fitted parameter. This function choice was made due to the observed almost linear dependency in log-log coordinates (figure 1 (Left)) but with the added constraint that $D_{\text{qat}}\leq D_{\text{total}}$. The optimal fit in our setup yields $a=6.7297$.
Optimal QAT Fraction Fit Results.
We fit the proposed equation directly using Huber loss (Huber, 1964) and gradient descent optimization. The approach achieves 0.091 MAE in fraction prediction across all model sizes and experiments. We also verify that the error remains low if we remove the largest tested model from the training and evaluate accuracy only on it. The results are displayed in figure 1 (Left). Optimal points lie close to the predicted optimal fraction. We can make the following high-level observations: the optimal QAT fraction grows faster with $D_{\text{total}}$ for lower bit widths, the optimal QAT fraction decreases with model size $N$ increase and fixed $D_{\text{total}}$. One limitation of the fit is that it is subject to the granularity of the selected experiments—the set of QAT fractions being tested. Also, as we fit only optimal points, we do not use a substantial amount of non-optimal data points, which also contain valuable information about loss behavior. One way to utilize all available data is to model the loss scaling law explicitly and infer optimal QAT fractions from it. We focus on this in section 4.2.
4.2 Loss Scaling Law
As described in section 2, Chen et al. (2025b) were able to fit a loss scaling law for QAT started from scratch ($D_{\text{fp}}=0$). We extend this idea by making the loss scaling law dependent not only on $D_{\text{total}}$ but also on $D_{\text{fp}}$, $D_{\text{qat}}$, and $B$, essentially modeling loss for different QAT fractions and bit-widths. However, we do not follow the same functional form as that proposed by Chen et al. (2025b). This is because in equation 2.1, the QAT penalty overtakes the Chinchilla loss term at some point with the growth of token count, which causes the limit of the whole expression to approach infinity as $D_{\text{total}}\to\infty$. This does not align with the expected loss decrease as token count grows and will hinder making any predictions from the scaling law in the future. We propose a loss model in the form of
[…]
where all lowercase Greek letters are fitted parameters and $S_{\text{qat}}=\frac{D_{\text{qat}}}{N\cdot\frac{B}{8}}$, $S_{\text{fp}}=\frac{D_{\text{fp}}}{N\cdot\frac{B}{8}}$ are the corresponding tokens-per-parameter-byte. This choice of $\delta(N,D_{\text{qat}},D_{\text{fp}},B)$ is motivated by the dependence of the optimal QAT fraction on tokens-per-parameter-byte as discussed in section 4.1; specific motivation for each term is described in equation 4.1.
Loss Scaling Law Fit Results.
We fit the proposed equation for 757 total QAT experiments directly using Huber loss (Huber, 1964) and gradient descent optimization—a setup consistent with that of Hoffmann et al. (2022b); Chen et al. (2025b). The results are highly dependent on initialization; therefore, we select the best fit out of many random initializations. We achieve similar fit quality across different bit-widths: $R^{2}=0.982$ for 1-bit QAT and $R^{2}=0.991$ for 6-bit QAT, where $R^{2}$ is the coefficient of determination. Full fit metrics are presented in table 1. We present the fitted formula and its visualizations in appendix 12 and plot the 3D loss scaling law surface for fixed model size in figure 4. Also, in appendix 14 we fit formulas independently for each bit-width $B$ as a baseline and verify that the unified formula achieves similar fit metrics. Additional scaling law fit notes are provided in appendix 13, and appendix 18 verifies that optimal QAT fraction prediction generalizes to larger model sizes (2.2B).
Table 1: Fit metrics for the loss scaling law. We report both the metrics of the loss fit and of the optimal QAT fraction prediction inferred from the loss scaling law. It is seen that the proposed formula in equation 4.1 provides a good fit of loss as well as of the optimal QAT fraction.
$\bm{B}$ MAE, loss fit $\bm{R^{2}}$, loss fit MAE, optimal QAT fraction fit
1 0.026 0.982 0.081
2 0.023 0.981 0.102
4 0.021 0.983 0.074
6 0.018 0.991 0.09
The fitted formulas are analytically sound: with an increase of either $D_{\text{fp}}$ or $D_{\text{qat}}$ while the other is fixed, the total loss decreases. Additionally, the proposed form effectively captures the optimal QAT fraction. From experimental results and the fitted loss function, we observe that low-bit QAT is more sensitive to the QAT fraction being optimal. Loss increase for sub-optimal QAT is higher for 1-bit than for 6-bit. Therefore, selecting the optimal QAT fraction is especially important in low-bit settings. To assess the generality of our formulation, we reproduce our findings with different pre-training and QAT hyperparameters and the SlimPajama (Soboleva et al., 2023) dataset in appendix 17.
Quantifying Sub-Optimality.
After obtaining loss scaling laws, we can analyze how the optimal QAT setup compares to a sub-optimal one. To do so, we calculate wasted token count—the number of tokens that are effectively wasted by a sub-optimal QAT fraction. Wasted token count is summarized in figure 4 for different bit widths and token counts; we use 10% QAT as the reference for the sub-optimal setup. Two factors influence the wasted token magnitude: closeness of 10% to the optimal QAT fraction and the overall flatness of the loss scaling law for high token counts. If the predicted loss is generally flat for some token count, then even high deviation from optimality will yield minor wasted token count. In the extreme case, for 1-bit QAT, the same loss can be achieved with just around 50% of compute if the optimal QAT fraction is used. This effect is still present for 2–4-bit QAT but becomes relatively small for 6-bit.
[…]
[width=]svg-inkscape/3dloss_svg-tex.pdf_tex
Figure 3: Visualization of fitted loss scaling law for 759M model, 1-bit QAT, and different $D_{\text{qat}},D_{\text{fp}}$. Orange lines represent constant $D_{\text{total}}=D_{\text{qat}}+D_{\text{fp}}$ levels, and stars represent loss minima for each such level. It is clearly seen that the loss structure yields an optimal QAT fraction for a specific $D_{\text{total}}$. The overall phenomenon is consistent with what was discussed in section 4.1.
[…]
[width=]svg-inkscape/subopt_compare_svg-tex.pdf_tex
Figure 4: Comparison of sub-optimal QAT setup with fixed 10% QAT fraction and optimal QAT setup for 1B parameter model. Wasted token count is the number of tokens effectively wasted by not utilizing an optimal QAT fraction setup. That is, if the wasted token count is $n\%$, then the same loss can be achieved with $(100-n)\%$ tokens and optimal QAT fraction. While results vary for different bit widths, the general relationship is similar, revealing high potential savings.
Note on Compute Budget/Token Budget Duality.
So far, we have considered token count to be identical to compute budget as compute scales linearly with training token count. However, one may argue that since QAT employs additional operations, its complexity is higher than FP training. As QAT overhead depends only on model size, it becomes negligible with sufficiently large batch size and sequence length (appendix 19). Still, in setups where QAT overhead is substantial, one can obtain compute-based optimal QAT fraction from the token-based approach by making a substitution $D_{\text{qat}}=\frac{1}{r}\cdot D_{\text{qat}}^{\prime}$, where $r>1$ is the QAT overhead factor in the specific setup and $D_{\text{qat}}^{\prime}$ is the overhead-aware QAT token count. This will make QAT steps “more expensive” from a loss minimization perspective. In this setup, $D_{\text{fp}}+D_{\text{qat}}^{\prime}=const$ represents not iso-token levels, but rather iso-flop levels. Therefore, the inferred optimal QAT fraction will be adjusted to account for the overhead.
5 When Does QAT Restore FP Model Accuracy?
An interesting question to analyze is how QAT loss relates to FP model loss for the same total token count. Specifically, a practically significant question is “when does QAT restore FP model accuracy?” To answer this question, we plot the difference in perplexities between QAT and FP models for each total token count; appendix 13 describes how we obtain the full-precision model loss scaling law. Figure 5 presents such a plot for models of two different sizes.
FP Accuracy Reproduction.
The practical observation that larger-parameter models can tolerate lower-bit QAT is clearly observed. A second perspective from which to consider figure 5 is that of optimal QAT bit width. Specifically, for a given total token count, there exists a minimum bit width that matches FP model loss. Therefore, there is no incentive to train higher-bit QAT, as this will not result in better accuracy but only in higher memory usage.
[…]
[width=]svg-inkscape/fp_qat_diff_plot_svg-tex.pdf_tex
Figure 5: Difference in perplexity between FP loss scaling law and QAT loss scaling law for two model sizes. For QAT, the loss corresponding to the optimal QAT fraction is used. Values below 0 correspond to QAT performing better than the FP model. It is clearly observed that the ability of QAT to match FP loss is greatly influenced by model size and token count. In particular, larger models are able to tolerate lower QAT precision for higher total token count budgets. Additional plot information is present in appendix 15.
6 Parameter–Precision Trade-Off
An interesting question to analyze is “for a fixed model memory requirement, how should one select QAT precision and parameter count?”. That is, to fit a specified memory constraint, one can choose high precision at the expense of lower parameter count or vice versa. This question is practically important as LLM inference is largely bottlenecked by memory bandwidth (Davies et al., 2025; Recasens et al., 2025; Dao et al., 2022). We can derive such optimality from the fitted loss scaling law.
[…]
[width=]svg-inkscape/mem_opt_qat_pred_svg-tex.pdf_tex
Figure 6: Optimal QAT bit width for different memory budgets and total training budgets. We use the loss corresponding to the optimal QAT fraction. For training FLOPs, we use the estimation $C\thicksim 6ND$. The white area corresponds to $D<N$, which is not practically important.
The results are presented in figure 6. It is clearly seen that for a fixed memory budget, optimal QAT precision decreases with training FLOP growth. This suggests that for achieving an optimal quantized model within some memory and training compute budget, one should select parameter count in advance accordingly. We believe this finding to be important for practitioners trying to achieve the best-accuracy model within memory constraints. Figure 6 is verified experimentally in appendix 16.
7 Cooldown & QAT Fusion
Section 4 revealed the importance of advance planning for QAT, accounting for the optimal fraction. This is possible only when one controls the entire pretraining process: both QAT and FP. In this context, it may be worth adjusting the training procedure to make QAT more efficient. Specifically, in this section we focus on modifications to learning rate scheduling techniques. Currently, a classic way of training models is to perform full FP training with learning rate cooldown, and then start QAT with learning rate re-warmup. We used such a setup for the scaling law in section 4 as it is universally adopted. However, a more optimal scheme may exist.
QAT & Learning Rate Cooldown Fusion.
Wen et al. (2024) show that re-initializing WSD from a post-cooldown checkpoint rather than from a constant stage yields better results. However, we believe the behavior might be different when resuming training from a checkpoint with QAT. We propose a novel idea for QAT: QAT & learning rate cooldown fusion. Motivated by the idea that learning rate cooldown performs low-magnitude adjustments to weights, we speculate that a substantial part of updates during learning rate cooldown gets destroyed by QAT initialization, which, in essence, discards high-precision information. Therefore, we analyze a setup where QAT is started directly from the learning rate constant stage and learning rate cooldown is performed jointly with QAT. A schematic representation of the two schemes is presented in figure 7.
[…]
[width=]svg-inkscape/cooldown_qat_fuse_svg-tex.pdf_tex
Figure 7: Comparison between two different QAT schemes. In both setups, the QAT fraction is 40%. Red-shaded areas indicate zones with lowered learning rate, which we expect to correspond to minor weight updates that get effectively ignored by QAT initialization. On the left, classic QAT scheme visualization: QAT follows fully completed FP training that ends with 20% (of FP training length) learning rate decay. For QAT, the learning rate follows a cosine shape with 5% re-warmup phase. On the right, the cooldown & QAT fusion scheme is displayed. QAT starts directly from the constant learning rate stage with small re-warmup, effectively resuming the FP learning rate scheduler as if QAT was not present at all. QAT ends with 20% cooldown (of total training length). As QAT follows the classic FP learning rate recipe with usual cooldown, we call this approach cooldown & QAT fusion.
We run experiments for different model sizes and 4-bit QAT with the described cooldown & QAT fusion scheme, taking experiments with the classic QAT scheme and optimal QAT fraction as baselines. The results are displayed in table 2. In addition to perplexity, we report loss change in “wasted tokens” units. This is the total token distance between corresponding loss points in the scaling law for an optimal QAT fraction. Such a metric is reported for better impact understanding, as small perplexity differences are harder to achieve with high overall token counts. We achieve improvements across all model sizes for 4- and 6-bit widths and all token counts. Results differ for 1- and 2-bit settings; we believe this is due to the large optimal QAT fraction, which makes the FP fraction small and, consequently, the impact of cooldown & QAT fusion lower. Full results are available in appendix 20. While perplexity differences may seem small, judging the difference from the perspective of token count difference is significant. This implies substantial improvements in terms of training cost.
Table 2: Accuracy comparison between the classic QAT scheme and the cooldown & QAT fusion training scheme. The loss difference is reported in “wasted tokens”—the difference in total token count between optimal QAT fraction loss points in the loss scaling law. Substantial improvements are noticeable across different model sizes and token counts.
Perplexity Wasted tokens
Unfused Fused Unfused total tokens, %
$\bm{B}$ Model size Total tokens
4 74 1.4T 16.26 16.25-0.06% 2.2%
163 901.3B 13.51 13.49-0.15% 9.2%
425 10.5B 16.3 16.02-1.72% 9.6%
31.8B 13.9 13.76-1.01% 10.4%
96.0B 12.62 12.54-0.63% 13.6%
816 281.9B 11.07 11.02-0.45% 13.2%
8 Conclusion
This work addresses a resource allocation problem in quantization-aware training: how to optimally divide compute between full-precision pretraining and quantization-aware training. Through extensive experiments across model sizes, compute budgets, and quantization bit widths, we challenge existing assumptions and provide practical guidelines for efficient QAT planning. Our key contributions are:
  • Discovery of Compute-Dependent Optimal QAT Fractions. Through extensive experiments across different model sizes, compute budgets, and QAT bit widths, we demonstrate that previous assumptions about optimal QAT allocation do not hold as compute budgets increase. Our findings reveal that the optimal QAT fraction is not a fixed percentage but rather increases with the total compute budget, specifically with the tokens-per-parameter-byte statistic. This challenges the previous conclusion that 10% is universally optimal for QAT length relative to total training length. We demonstrate that using suboptimal QAT fractions can result in substantial compute waste, with extreme cases showing that the same loss can be achieved with just around 50% of the compute when optimal QAT fractions are used, particularly for low-bit quantization scenarios.
  • Comprehensive Loss Scaling Law. We derive a comprehensive loss scaling law that models the final expected loss of the full-precision and quantization-aware training pipeline as a function of QAT bit width, model parameter count, and token counts for both training phases. This scaling law not only captures the optimal QAT fraction phenomenon but also enables prediction of final model loss across different QAT/FP allocation strategies. From the scaling law, we infer which QAT bit width is optimal under a given memory constraint and how QAT accuracy compares to FP model accuracy.
  • Cooldown and QAT Fusion Technique. We introduce a novel approach that performs learning rate decay jointly with quantization-aware training, eliminating redundant full-precision updates and achieving significant compute savings.
Limitations.
While we performed experiments with different datasets and hyperparameters, our work still focuses on a specific LLM architecture, and exact results may differ for different model types. However, we expect the overall observed phenomena to be consistent across different architectures.
Future Work.
We identify several research directions worth exploring. First, the relationship between the optimal QAT fraction and pretraining precision remains unknown. This direction is especially interesting with the emergence of 8-bit floating-point training (Peng et al., 2023) and even 4-bit training (Zhou et al., 2025b; Wang et al., 2025). Second, we are interested in how the observed phenomena are preserved across different training stages. Specifically, how does the optimal QAT fraction change when the full-precision training stage incorporates additional stages such as Supervised Fine-Tuning (SFT) (Lee, 2024), Reinforcement Learning (RL) (Rafailov et al., 2023; Chen et al., 2023), or multimodal training? We speculate on these questions in appendix 23.
Reproducibility Statement.
We report exact hyperparameters and training approaches used in section 3. Additional experimental information that should facilitate reproduction is summarized in appendix 10, 11, and 21.
Ethics Statement: LLM Use Disclosure.
LLMs such as Anthropic (2025); Mistral AI (2025) were used in the preparation of this paper exclusively for improving grammar and wording.